OpenStudy (anonymous):

The 14 students in Mateo's computer programming class are using the computer lab. Some computers are shared by pairs of students, and 2 students work alone. Which equation could you use to determine how many computers are shared in the computer lab? Let c be the number of shared computers. (Answer options in reply bar)

OpenStudy (davidusa):

i dont know sorry

OpenStudy (davidusa):

i think 6 computers

OpenStudy (anonymous):

A. \[\frac{ c }{ 2 } - 2 = 14 \] B. \[2c - 2 = 14\] C. \[2c + 2 = 14\] D. \[\frac{ c }{ 2 } + 2 = 14\]

OpenStudy (davidusa):

i think it is d.

OpenStudy (anonymous):

Thank you c:

OpenStudy (davidusa):

thanks for medals. you didnt message me!!!

OpenStudy (anonymous):

Lol sorry I forgot xD

OpenStudy (davidusa):

ok